Solve the equation by expressing each side as a power of the same base and then equating exponents.
e^x+5=1/e^6

Answer by stanbon(75887)

Solve the equation by expressing each side as a power of the same base and then equating exponents.
e^(x+5) = 1/e^6
---
e^(x+5) = e^(-6)
---
Equate the exponents:
x+5 = -6
x = -11
===================
Cheers,
Stan H.